OpenStudy (anonymous):

The sum of three consecutive odd integers is –27. What are the numbers? A. {–5, –9, –13} B. {–10, –9, –8} C. {–7, –9, –11} D. {–11, –13, –15}

OpenStudy (jojo4eva):

i believe the answer would be C

OpenStudy (anonymous):

Im gonna go with B since they are consecutive order

OpenStudy (jojo4eva):

ya it is B cuz its going from least to greatest so ya its B

OpenStudy (texaschic101):

x = 1st integer x + 2 = 2nd integer x + 4 = 3rd integer x + x + 2 + x + 4 = -27 3x + 6 = -27 3x = -27 - 6 3x = -33 x = -11 x + 2 = -11 + 2 = -9 x + 4 = -11 + 4 = - 7 It is C